The following questions refer to an electron in a hydrogen atom relaxing from the n=3 energy level to the n=1 energy level.

Use the energy (-1.936×10-18 J) to calculate the energy for this transition in kilojoules per mole of photons.

Round your answer to the correct number of significant figures.
Answer #1

Energy for single photon = -1.936×10-18 J

= -1.936×10-21 KJ (1KJ = 1000 J)

Now, 1 mole contain 6.023×1023 photons

Thus, energy of 1 mole photon= energy of 6.023×1023 photons

= -6.023×1023×1.936×10-21 KJ/mol

= -11.67 ×102 KJ/mol

= - 1167 KJ/mol